OpenStudy (anonymous):

what is the dlope of the points 2,14 and 0,-12

OpenStudy (anonymous):

i mean slope

Directrix (directrix):

[14 - (-12)] / (2 - 0) = 26/2 = 13

Directrix (directrix):

Final Answer: 13